OpenStudy (anonymous):

Use factoring and the zero product property to solve equation. 2x^2 -9x-5=0

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

\(2\times -5=10\\ -9-1=10\)

OpenStudy (mathmath333):

do u know how to factor ?

OpenStudy (anonymous):

I think @mathmath333

OpenStudy (mathmath333):

ok can u show ur work

OpenStudy (anonymous):

What do i factor

OpenStudy (mathmath333):

lol the equation in the question

OpenStudy (anonymous):

oh lol no im not sure how to do it

OpenStudy (mathmath333):

but u can try , ain't u?

OpenStudy (anonymous):

Im just really having a total brain fart right now

OpenStudy (mathmath333):

so u dont know how to factor

OpenStudy (anonymous):

no

OpenStudy (mathmath333):

ok no problem

OpenStudy (anonymous):

how do we do it

OpenStudy (mathmath333):

\(\Large\ \begin{align} \color{black}{2x^2 -9x-5=0\\ 2x^2 -10x+x-5=0\\ 2x(x -5)+(x-5)=0\\ (2x+1)(x -5)=0\\~\\ \text{zero product rule}\\~\\ (2x+1)=0~~or~~(x -5)=0\\ x=\dfrac{-1}{2}~~or~~x =5}\end{align}\)